OpenStudy (anonymous):

x−1=74x−1+40+y−2+180 x=75 and y=68? Need steps am confused

OpenStudy (anonymous):

do you have to find the value of a specific variable?

OpenStudy (anonymous):

@mmk1

OpenStudy (anonymous):

yes x and y x=75 y=68 but confused with steps

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

180=z+13+Y SAME TYPE-NEED TO SOLVE FOR Z

OpenStudy (anonymous):

x−1=74x−1+40+y−2+180 First things first....(combine like terms) x−1=74x−139+y -74x-1=-139 + y +1 1 ---------------- -74x=-138 + y What did I do wrong?????......... Let's see show me how you got X AND Y

OpenStudy (anonymous):

The second one is: 180=z+13+Y -13 -13 ------------ 167=z + Y - Y - Y ---------- 167 - Y = z or z = -Y + 167

OpenStudy (anonymous):

That's how u solve for Z

OpenStudy (anonymous):

OK THANK YOU!

OpenStudy (anonymous):

np how bout the first one?

OpenStudy (anonymous):

BACK TO THE FIRST ONE -74x=-138+y -74(75)=-138+68 -5550=-70???

OpenStudy (anonymous):

at first like, you solve it and show me r work

OpenStudy (anonymous):

x-1=74x-1+40+y-2+18 x=74x+37+18 x=74x+55 I am stuck

OpenStudy (anonymous):

@phi @ganeshie8 can u plz help mmk1 with this

OpenStudy (phi):

if you have 2 unknowns then (to solve) you need 2 separate equations

OpenStudy (anonymous):

Your equation (if the above equation as mentioned in the question is correct) will become like this => 73x + 218 + y = 0 Now 2 things can be inferred : 1) We need another equation to solve for values of x and y. 2) The nature for the ans. for x and y can never be simultaneously positive. One of them will surely be a negative no.